### v1

# Learning to Guide Human Decision Makers with Vision-Language Models

Abstract
There is increasing interest in developing AIs for assisting human decision making in high-stakes tasks, such as medical diagnosis, for the purpose of improving decision quality and reducing cognitive strain. Mainstream approaches team up an expert with a machine learning model to which safer decisions are offloaded, thus letting the former focus on cases that demand their attention. This separation of responsibilities setup, however, is inadequate for high-stakes scenarios. On the one hand, the expert may end up over-relying on the machine’s decisions due to anchoring bias, thus losing the human oversight that is increasingly being required by regulatory agencies to ensure trustworthy AI. On the other hand, the expert is left entirely unassisted on the (typically hardest) decisions on which the model abstained. As a remedy, we introduce learning to guide (LTG), an alternative framework in which – rather than taking control from the human expert – the machine provides guidance useful for decision making, and the human is entirely responsible for coming up with a decision. In order to ensure guidance is interpretable and task-specific, we develop slog, an approach for turning any vision-language model into a capable generator of textual guidance by leveraging a modicum of human feedback. Our empirical evaluation highlights the promise of slog on a challenging, real-world medical diagnosis task.
1 Introduction
High-stakes applications in healthcare, criminal justice and policy making can substantially benefit from the introduction of AI technology, yet full automation in these scenarios is not desirable, due to ethical, safety and legal concerns, if not explicitly forbidden by law (Government of Canada, 2019; European Commission, 2021). For these reasons, human-AI or Hybrid Decision Making (HDM) is becoming increasingly popular to tackle high-stakes tasks. HDM algorithms pair a human decision maker with an AI agent – often a machine learning model – capable of providing support, with the goals of improving decision quality and lowering cognitive effort.
Most current approaches to HDM follow a principle of separation of responsibilities, in the sense that they work by routing novel inputs to exactly one of the two agents – either the human or the AI – who is then responsible for coming up with a decision. Specifically, in existing approaches (Madras et al., 2018; Mozannar and Sontag, 2020; Keswani et al., 2022; Verma and Nalisnick, 2022; Liu et al., 2022; Wilder et al., 2021; De et al., 2020; Raghu et al., 2019; Okati et al., 2021), the AI first assesses whether an input can be handled in autonomy – e.g., it is low-risk or can be addressed with confidence – and defers to a human partner otherwise. These algorithms are beneficial in that they enable the human to focus on those cases that (according to the machine) most require their attention.
We argue that this setup is suboptimal and potentially unsafe. It is suboptimal because, whenever the machine opts for deferral, the human is left resolving hard cases completely unassisted (as in Fig. 1, right). At the same time, it is unsafe, because humans are affected by anchoring bias (Rastogi et al., 2022; Eigner and Händler, 2024), a phenomenon whereby decision makers tend to blindly rely on an initial impression (the anchor) and refrain from exploring alternative hypotheses. When the anchor is provided by an algorithm, the bias is amplified as humans tend to over-trust the machine’s decisions when available and ignore their own opinions, a phenomenon called automation bias (Cummings, 2012) (Fig. 1, middle). This effectively undermines human oversight over algorithmic decisions, which is increasingly being required by governments around the world to regulate the use of AI in high-stakes applications (Green, 2022).
Figure 1: Left: Existing HDM approaches employ a deferral function $d(\bm{\mathrm{x}})$ to partition the input space $\mathcal{X}$ into $\mathcal{H}$ and $\mathcal{M}$. Middle: A predictor $f(\bm{\mathrm{x}})$ handles those inputs falling in $\mathcal{M}$ (blue arrow). Because of anchoring bias, the human expert may end up blindly trusting its (possibly poor) decisions $y_{m}$. Right: The human, on the other hand, is left completely unassisted for those (possibly hard) decisions falling in $\mathcal{H}$, increasing the chance of mistakes in the human’s decisions $y_{h}$ (green arrow).
As a remedy, we propose learning to guide (LTG), an alternative algorithmic setup that side-steps these issues. In LTG the machine is trained to supply its human partner with interpretable guidance highlighting those aspects of the input that are useful for coming up with a high-quality decision. For instance, in pathology prediction, the guidance could highlight those pathology present in an input X-ray scan that are indicative of possible diagnoses. In LTG, by construction, all decisions are taken by the human expert – thus preventing automation bias – but facilitated by accompanying machine guidance.
We showcase LTG on medical decision making focusing on guidance formulated in natural language. To this end, we introduce slog (Surrogate-based Learning tO Guide), an algorithm for turning large vision-language models (VLMs) (Radford et al., 2021; Yan and Pei, 2022; Sharma et al., 2021) into high-quality guidance generators. In a nutshell, slog takes a VLM pre-trained for caption generation and fine-tunes it using feedback about the quality of downstream human decisions inferred from generated guidance. slog keeps annotation costs under control by training a surrogate model that predicts downstream decision quality on a modest amount of feedback, and then using it to fine-tune the VLM in an end-to-end fashion. Our experiments on a challenging medical diagnosis task indicate that VLMs fine-tuned with slog output interpretable task-specific guidance that can be used to infer high-quality decisions.
Contributions.
Summarizing, we:
  • Identify serious flaws with existing HDM algorithms that compromise their applicability to high-stakes tasks.
  • Introduce learning to guide (LTG), a novel approach for assisting human decision makers that ensures they are always in the loop.
  • Develop slog, an LTG approach tailored for natural language guidance that can convert large VLMs into interpretable, task-specific guidance generators.
  • Showcase the promise of slog on a challenging medical diagnosis task.
2 Hybrid Decision Making
We are concerned with decision tasks that, due to safety concerns, cannot be fully automated, such as medical diagnosis. Specifically, we focus on classification problems, with inputs $\bm{\mathrm{x}}\in\mathbb{R}^{d}$ and categorical or multi-label decisions $y$ [^1] While we focus on classification tasks, our remarks apply to other prediction problems, e.g., regression (De et al., 2020).
Research on Hybrid Decision Making (HDM) investigates expert AI assistants capable of assisting human experts in such tasks. Considering the AI assistant and the human expert have different abilities, expertise, and biases, the central question of HDM is how to best integrate them.
2.1 HDM by Separation of Responsibilities
Existing HDM strategies solve this problem by following a principle of separation of responsibilities: any given instance $\bm{\mathrm{x}}$ is assigned to exactly one of the two agents, who is then in charge of decision making, as shown in Fig. 1. Specifically, they implement a classifier $f:\bm{\mathrm{x}}\mapsto\hat{y}$, playing the role of an AI agent, as well as a deferral policy $d:\mathbb{R}^{d}\to\{{\tt machine},{\tt human}\}$ that partitions the input domain $\mathcal{X}$ into two disjoint subsets, $\mathcal{M}$ and $\mathcal{H}$. Novel inputs $\bm{\mathrm{x}}$ falling in the former are handled by $f$ and those falling in the latter are handled by the human expert. This setup is known under a variety of names, including learning to defer (Madras et al., 2018; Mozannar and Sontag, 2020), learning under algorithmic triage (Raghu et al., 2019; Okati et al., 2021), learning under human assistance (De et al., 2020, 2021), and learning to complement (Wilder et al., 2021; Bansal et al., 2021).
Approaches differ in how they partition the space of inputs $\mathcal{X}$. Earlier methods build on prediction with a reject option (Cortes et al., 2016), in which the deferral policy $d$ observes all incoming instances $\bm{\mathrm{x}}$ and offloads those about which the predictor $f$ is unsure (based on, e.g., predictive variance) (Raghu et al., 2019). Since $f$ is fixed, the partition is static and depends only the self-assessed uncertainty of the predictor. Assuming the latter is sufficiently well calibrated (Kendall and Gal, 2017), this strategy can perform well in practice (Liu et al., 2022). The main drawback with this setup is that the partitioning accounts for the machine’s performance only, neglecting the human’s expertise and biases. (Madras et al., 2018) improve on this by learning the deferral policy $d$ so that it optimizes some (decision theoretic) measure of joint team performance, and as such they explicitly take the quality of human decisions into account. Follow-up works (De et al., 2020, 2021; Wilder et al., 2021) go one step further and train the deferral policy $d$ and the predictor $f$ jointly, so as to adapt one to the other.
This setup has been extended to incremental (Keswani et al., 2021) and sequential (Joshi et al., 2021) decision making, and to bandit feedback (Gao et al., 2021). Theoretical studies have analyzed the consistency (Mozannar and Sontag, 2020) and calibration (Verma and Nalisnick, 2022) of the HDM pipeline and the structure of optimal deferral policies (Okati et al., 2021).
2.2 Issues with Separation of Responsibilities
At a high level, the desiderata that an HDM strategy should satisfy in the context of high-stakes decision making problems are the following:
  • D1.
    Complementarity. The human-machine team should leverage the complementary capabilities of each agent to obtain better decisions on average, or equally good decisions at a lower cognitive cost, than each agent individually.
  • D2.
    Synergy. The human-machine team should combine the contributions of each agent to obtain better individual decisions, or equally good decisions at a lower cognitive cost, than each agent individually.
  • D3.
    Reliability. The human-machine team should produce decisions that are more reliable than those made by each agent individually.
Existing HDM approaches aim specifically at enabling complementarity (D1), for two reasons. First, the main benefit of offloading decisions to an AI is that of lowering the human’s cognitive effort. Second, depending on the relative performance of the predictor on inputs falling in $\mathcal{M}$ compared to the expert, they can also improve the quality of the team’s decisions (on average across inputs, not necessarily for all inputs). In fact, under suitable conditions, learning to defer can provably do so (Donahue et al., 2022).
However, approaches complying with separation of responsibilities completely overlook synergy (D2). When the machine outputs a decision, the human is tempted to simply stick to it, thus over-relying on the machine’s decisions, because of the previously mentioned anchoring (Rastogi et al., 2022) and automation (Cummings, 2012) biases. Conversely, whenever the machine opts for deferral, the human is left resolving hard cases completely unassisted. This also compromises reliability (D3), which is key in high-stakes applications, thus hindering the applicability of HDM.
Figure 2: The slog approach to learning to guide. Top: Given an input $\bm{\mathrm{x}}$, slog uses a VLM $\gamma$ to output textual guidance $g$ in support of human decision making. Here, $q$ indicates the quality of the human’s downstream decision. Middle: The surrogate $\sigma_{\mathrm{quality}}$ estimates the quality of the downstream decisions and it is trained using a modicum on annotated guidance-quality pairs. Bottom: Given a trained surrogate $\sigma_{\mathrm{quality}}$, slog fine-tunes the VLM to output guidance $g$ achieving high (estimated) decision quality.
3 Beyond Separation of Responsibilities with Learning to Guide
We propose learning to guide (LTG), a novel HDM framework that addresses the shortcomings of existing strategies by foregoing separation of responsibilities. In a nutshell, LTG aims to learn a guidance generator $\gamma(\bm{\mathrm{x}})$, implemented as a machine learning model, that given an input $\bm{\mathrm{x}}$, outputs guidance $g$ that is useful for assisting human decision making on that input. In medical diagnosis, for instance, given a chest X-ray scan $\bm{\mathrm{x}}$, the guidance $g$ might describe pathology visible in the image that are useful for identifying pathologies and prescribe treatment, as shown in Fig. 2 (top). Critically, and in stark contrast with existing HDM approaches, in LTG the machine does not replace the human: the final decision is always taken by the human partner, in collaboration with the AI. This means the decision maker is always in the loop and responsible for the final decision.
3.1 Desiderata for Guidance
In order for generated guidance to support human decision making, it should satisfy the following properties:
  • D4.
    Interpretability. It should be understandable for the human expert at hand.
  • D5.
    Informativeness. It should be informative for the decision at hand.
If these are satisfied, then guidance can be used by human experts to address a specific downstream decision making task. Note that, satisfying these desiderata encourages satisfaction of D1D3. In fact, if guidance is interpretable (D4) and extracts decision-relevant elements from the input (D5), it should help the human in taking accurate decisions on individual instances (D2) and as a consequence improving the average quality of the decisions being made (D1). Additionally, interpretability helps the human in judging the quality of the guidance received, and thus evaluate the reliability of the overall decision (D3).
3.2 Learning to Guide for VLMs
In this paper we focus on textual guidance expressed in natural language as a mean to enable interpretability (D4). Motivated by their state-of-the-art performance in text generation tasks (Wei et al., 2022) and by their promise in pathological report generation (Shamshad et al., 2023; Chen et al., 2020, 2021; Hou et al., 2021; Kayser et al., 2022; Yunxiang et al., 2023; Bazi et al., 2023; Drozdov et al., 2020; Yan and Pei, 2022), we propose to leverage vision-language models (VLM) to generate guidance.
Off-the-shelf VLMs are not conceived for generating guidance for specific decision making tasks, and thus violate informativeness (D5). Clearly, a perfectly accurate medical report is also an optimal guidance for follow-up decisions, but generating highly accurate reports requires massive amounts of supervision, and reports generated by specialized VLMs are far from perfect (Shamshad et al., 2023).
The question is then how to convert such models into high-quality guidance generators. Focusing on (medical) decision making from image data, we address this problem by introducing slog, a novel approach for turning vision-language models into guidance generators using human feedback designed to comply with D1D5. The rationale behind slog is to encourage VLMs to focus on accurately reporting those aspects of the input image that are most relevant for the follow-up decisions, possibly overlooking less important details. Next, we briefly discuss how slog uses annotations and then proceed to outline the main algorithm.
3.3 Estimating Downstream Decision Quality
Optimizing guidance for synergy (D2) requires knowing the quality of downstream decisions taken by a human expert supplied with the guidance itself. slog assumes access to quality ratings $\bm{\mathrm{q}}\in[0,1]^{d}$, where each $q_{i}$ encodes the quality of a downstream decision. For instance, if the expert has to determine the state of two conditions (e.g., pneumonia and fracture), then $d=2$. Quality ratings for expert decisions can be obtained by comparing these against a gold standard (using, e.g., decision accuracy) or by consulting a second expert (using, e.g., a star rating system).
Clearly, there is a tension between the number of annotations necessary for fine-tuning a VLM and the cost of eliciting such annotations. slog addresses this issue by training a surrogate model $\sigma_{\mathrm{quality}}:(\bm{\mathrm{x}},\bm{\mathrm{z}})\mapsto\widehat{\bm{\mathrm{q}}}$ using a modicum of annotated quality ratings, and using it to estimate the quality of guidance $g$ generated by the VLM during fine-tuning. In practice, slog fits the surrogate on a training set $\mathcal{D}_{\mathrm{surr}}=\{(\bm{\mathrm{x}}_{i},\bm{\mathrm{z}}_{i},q_{i})\}$, where $\bm{\mathrm{x}}_{i}$ is an input image, $\bm{\mathrm{z}}_{I}$ is the embeddings of the VLM’s guidance $g_{i}$ for that input, and $q_{i}$ is the quality of that guidance, by minimizing an average cross-entropy loss of the form:
$\frac{1}{|\mathcal{D}_{\mathrm{surr}}|}\sum_{(\bm{\mathrm{x}},\bm{\mathrm{z}},\bm{\mathrm{q}})\in\mathcal{D}_{\mathrm{surr}}}\ \frac{1}{d}\sum_{i=1}^{d}\mathsf{CE}(q_{i},\sigma_{\mathrm{quality}}(\bm{\mathrm{x}},\bm{\mathrm{z}})_{i})$ (1)
3.4 The slog Loop
In essence, slog takes a pre-trained VLM caption generator $\gamma$ and fine-tunes it for a number of rounds $T$. Let $\mathcal{D}_{\mathrm{train}}$ be a data set of image-caption pairs (for instance, a subset of the data that $\gamma$ was trained on) and $\mathcal{D}_{\mathrm{tune}}$ a larger set of unlabeled images from the target decision making task. In each round $t=1,\ldots,T$, slog samples a batch $\{\bm{\mathrm{x}}_{1},\ldots,\bm{\mathrm{x}}_{B}\}$ from $\mathcal{D}_{\mathrm{tune}}$ uniformly at random with replacement, and computes guidance $g_{i}^{t}=\gamma(\bm{\mathrm{x}}_{i})$ and embeddings $\bm{\mathrm{z}}_{i}^{t}$ for each input. Then, it evaluates the quality of the generated guidance using the frozen surrogate $\sigma_{\mathrm{quality}}$ and fine-tunes the VLM $\gamma$ by minimizing an augmented loss of the form:
$\mathsf{CE}(\gamma,\mathcal{D}_{\mathrm{train}})-\frac{\lambda}{|\mathcal{D}_{\mathrm{tune}}|}\sum_{(\bm{\mathrm{x}},\bm{\mathrm{z}})\in\mathcal{D}_{\mathrm{tune}}}\frac{1}{d}\sum_{i=1}^{d}\sigma_{\mathrm{quality}}(\bm{\mathrm{x}},\bm{\mathrm{z}})_{i}$ (2)
for a given number of epochs. Eq. 2 trades off text generation performance on the training set $\mathcal{D}_{\mathrm{train}}$ – so as to discourage catastrophic forgetting – and estimated quality of downstream decisions on the fine-tuning set $\mathcal{D}_{\mathrm{tune}}$. Here, $\lambda>0$ is a hyper-parameter. Fine-tuning then amounts to applying gradient descent to batches comprising training and fine-tuning examples in equal proportions. Once done, the slog loop then repeats. As long as the surrogate generalizes the quality rating annotations, the VLM gradually learns to output text that work well as guidance tailored for the target decision task.
3.5 Benefits and Limitations
In stark contrast with existing HDM strategies, slog ensures that the human receives guidance useful for decision making while keeping them in the loop. While LTG is more cognitively demanding than existing HDM approaches, these cannot be safely employed in high-stakes applications, where there is little room for mistakes and humans have to be in control at all times (Zhang et al., 2020), as increasingly prescribed by legal frameworks (Government of Canada, 2019; European Commission, 2021). LTG and slog are designed explicitly for supporting HDM in these cases. slog is reminiscent of mainstream approaches to LLM alignment, such as reinforcement learning with human feedback (RLHF) (Ziegler et al., 2020; Ouyang et al., 2022), but differs from them in aims and technology. While RLHF strives to improve factuality and reduce harmfulness of generated content (Ouyang et al., 2022), ignoring the decisions these impinge on, slog specifically aims at improving quality of downstream human decisions for a specific decision making task. At the same time, slog foregoes reinforcement learning approaches (Schulman et al., 2017) in favor of a simpler and more direct end-to-end fine-tuning strategy.
One limitation of slog is that the performance of the guidance generator hinges on that of the surrogate, which in turn relies on the amount of quality ratings available for training. In Section 4 we show that, however, a limited amount of quality annotations are sufficient for slog to substantially improve generated guidance. Another limitation of slog is that it currently assumes quality ratings are readily available, which is not always the case. One option is then to integrate slog with active learning strategies (Settles, 2012; Herde et al., 2021) to acquire informative quality ratings whenever needed. Doing so is however outside the scope of this paper and left to future work. Finally, the guidance output by VLMs may suffer from hallucination, that is, it may contain untrue statements. However, slog directly maximizes factuality of guidance on the fine-tuning set $\mathcal{D}_{\mathrm{tune}}$, meaning that a simple way of reducing the chance of non-factual statements is to employ a larger fine-tuning set. This is relatively cheap to do, as no annotations are required. Moreover, large language models can be surprisingly well-calibrated (Kadavath et al., 2022), meaning that generated guidance can be filtered based on the VLM’s own uncertainty estimates to prevent over-reliance (Eigner and Händler, 2024) and avoid low-quality decisions (Zhang et al., 2020).
Figure 3: An example of a radiography along with its corresponding medical report consisting of ‘findings’ and ‘impression’.
4 Empirical Analysis
In this section, we answer empirically the following research questions:
  • Q1.
    Does slog improve the informativeness of generated guidance despite relying on a surrogate model?
  • Q2.
    Does slog improve the quality of the decisions being made using its guidance?
Table 1: Comparison between the splits of the original data and our filtered data.
Original split Our split
Components Train Val Test Train Val Test
Reports 222,758 1,808 3,269 125,417 991 1,624
Images 368,960 2,991 5,159 232,855 1,837 2,872
Data set.
We evaluate slog on the Mimic-CXR-IV data set (Johnson et al., 2019), one of the largest publicly available medical decision data sets, consisting of $227,835$ radiology reports and $377,110$ chest X-ray scans. We focus on the findings and the impression sections of the reports. As shown in Fig. 3, the findings are text-based descriptions of what can be observed in the scan, and constitute the basis on top of which the expert forms their impression, i.e., their initial opinion about the potential pathologies of a patient. We discarded examples where either findings or impressions were not available, resulting in the training, validation and test splits presented in Table 1.
Vision-language Models.
We consider the R2Gen (Chen et al., 2020) vision-language model, a memory-driven transformer specifically designed for pathological report generation from chest X-ray images. The R2Gen architecture builds on the observation that similar radiographs may correspond to reports sharing similar patterns. To exploit this, it employs a pre-trained CNN model to extract patch features and encoding these into hidden states with an encoder. A decoder then maps the hidden states into words at each time point with the help of a relational memory and memory-driven conditional layer normalization. The relational memory component allows the transformer to store and repurpose shared patterns and thus generate more coherent reports (Chen et al., 2020).
The decision-making task.
Our study focuses on a critical step of the medical decision process: making the right diagnosis [^2] While strictly speaking medical diagnosis is not a decision making task, an incorrect diagnosis inevitably leads to poor decisions.. The task consists in diagnosing pathologies (see Table 3 for the list of $14$ possible pathologies) from X-ray images. In our experiments, we pre-train a R2Gen VLM to predict findings from images, and then fine-tune it with slog to automatically generate higher-quality guidance for a human decision maker. Given that the impression is the opinion that the expert forms about potential pathologies visible in the image, we fine-tune our VLM to produce textual guidance that – once interpreted by a human expert – leads to the same diagnosis entailed by the impression.
Simulating the human expert.
slog can do so provided it has access to quality ratings of downstream human decisions. We simulate human decisions using a machine learning model $\sigma_{\mathrm{human}}$, for reproducibility. Specifically, $\sigma_{\mathrm{human}}$ takes a scan $\bm{\mathrm{x}}$ and a corresponding VLM-generated report and diagnoses the $14$ candidate pathologies using three classes: definitely present (positive), definitely absent (negative), and unclear (ambiguous). Following (Lovelace and Mortazavi, 2020a), we implement $\sigma_{\mathrm{human}}$ as a bidirectional LSTM augmented with attention and train it on ground-truth labels obtained by applying the CheXpert (Irvin et al., 2019) automated annotation tool to the ground-truth impressions. In order to emulate a setting with sparse human supervision, we assume ground-truth labels are available for 10% of the training data only. This ground-truth dataset $\mathcal{D}_{\mathrm{surr}}$ is used for training both the model simulating human decisions $\sigma_{\mathrm{human}}$ and the quality surrogate model $\sigma_{\mathrm{quality}}$ estimating the quality of these decisions. We rely on a stratified sampling procedure to select $\mathcal{D}_{\mathrm{surr}}$ so as to maintain a reasonable coverage of the different classes. First, $\sigma_{\mathrm{human}}$ is trained on $\mathcal{D}_{\mathrm{surr}}$ to output diagnosis given ground-truth findings (are these are the only ones for which we know the corresponding human diagnosis). Once trained, we use $\sigma_{\mathrm{human}}$ to produce quality ratings by computing the correctness of its predictions over $\mathcal{D}_{\mathrm{surr}}$. To avoid biasing the quality rating supervision by computing it on training instances, we run $k$-fold cross validation on $\mathcal{D}_{\mathrm{surr}}$ and collect quality ratings from the $k$ validation folds. For each validation fold, we compute decisions using both VLM-generated guidance and ground-truth text, so as to provide examples of both predicted and ground-truth guidance to train the quality surrogate model.
Quality surrogate model.
As explained in Section 3.3, the surrogate $\sigma_{\mathrm{quality}}$ should estimate the quality of human decisions when fed with the VLM guidance. In our experiments, human decisions are proxied with the $14$ labels output by $\sigma_{\mathrm{human}}$. The surrogate is thus trained to predict the correctness of each of the $14$ predictions made by $\sigma_{\mathrm{human}}$. As for the machine model of the human $\sigma_{\mathrm{human}}$, the surrogate is also implemented as an attention-augmented LSTM and trained to minimize the average cross entropy loss on the ground-truth dataset described in the previous paragraph.
Overall pipeline.
Summarizing, our experimental pipeline is laid out as follows. In a first step, a R2Gen VLM is pre-trained on the training split $\mathcal{D}_{\mathrm{train}}$ to generate findings. The VLM is then applied to the decision ground-truth dataset $\mathcal{D}_{\mathrm{surr}}$ ($10\%$ of the training split) and the generated guidance is fed to $\sigma_{\mathrm{human}}$ to obtain (simulated) human decisions and corresponding quality annotations. This information is then used to fit the quality surrogate model $\sigma_{\mathrm{quality}}$. Finally, the VLM is fine-tuned on the entire training split according to Eq. 2 for $10$ epochs and evaluated on the test data.
Table 2: slog boosts estimated quality of generated guidance without compromising text quality. The results show that slog substantially improves estimated guidance quality as measured by the surrogate model ($\sigma_{\mathrm{quality}}$) without affecting text quality as measured by Bleu scores over ground-truth caption data.
[…]
R2Gen (pretrained) 0.36 0.22 0.15 0.11 47.5 0.20
R2Gen (fine-tuned) 0.37 0.23 0.15 0.11 48.9 0.31
slog 0.37 0.23 0.16 0.11 48.9 4.26
Figure 4: A qualitative example of the improvement of the guidance from slog with respect to the competitors. Green text indicates sentences that are (approximately) shared between the ground truth, slog and at least one of the competitors. Blue (resp. magenta) text indicates sentences shared between the ground truth findings (resp. impression) and slog, but missed by the competitors. No ground-truth sentences are shared between ground-truth text and competitors but missed by slog in this example.
Q1: slog improves informativeness on the test set without compromising BLEU score.
A potential issue with using a surrogate model as a proxy of decision quality is that the fine-tuned VLM might end up overfitting the surrogate and produce guidance that, while seemingly informative, is unrelated to the actual input scan. slog prevents this by complementing estimated guidance quality as computed by the surrogate model with guidance appropriateness for the input image as measured by cross entropy over a training set of findings. Table 2 confirms the effectiveness of this strategy. slog substantially improves estimated guidance quality (second term in Eq. 2) without compromising text quality, as measured in terms of BLEU score, over test examples. For the sake of fairness, we compared slog (with $\lambda=10$) with both the pre-trained R2Gen model (before the fine-tuning stage), and the R2Gen model fine-tuned for the same number of epochs as slog, but with caption-level supervision only (i.e., setting $\lambda=0$ in Eq. 2). Fig. 4 shows a qualitative example of the improvement in guidance of slog with respect to the competitors. First, slog’s guidance retains all pieces of text that any of the other approach shares with the ground-truth text (green text). On top of this, slog retrieves additional chunks of text that are shared with ground truth findings (blue text) and impression (magenta text), even if the latter are never explicitly included as training supervision, confirming the effectiveness of the quality surrogate in encouraging the generation of relevant guidance for the diagnosis. [^3] Notice that while the slog guidance text is longer than the one of the competitors in this example, with various chunks of text which are not obviously connected to ground-truth ones, its overall quality is still much higher (0.86 $F_{1}$, as compared with 0.4 and 0.5 for R2Gen and R2Gen fine-tuned respectively).
Q2: slog improves quality of decisions.
Table 3 shows the results in terms of decision quality, as measured by the $F_{1}$ score of the positive label for all 14 classes (multi-label prediction). Results clearly indicate the effectiveness of the slog guidance in improving decision quality, despite the little supervision it received. slog outperforms the competitors in 9 out of 14 classes. All methods fail to identify any positive occurrence for three particularly under-represented classes (Pneumonia, Pleural Other, Fracture), while slog slightly underperforms with respect to pre-trained R2Gen in 2 classes only. It is worth noticing that slog is especially effective in improving recall without affecting precision on average, as shown by the two bottom lines reporting results averaged over classes (macro) and instances (micro) respectively.
Table 3: slog boosts quality of downstream decisions. Results show per-class, macro and micro averaged precision, recall and $F_{1}$. Best $F_{1}$ results are boldfaced. slog outperforms competitors on 9 out of 14 classes, is on-par in 3 (where no method manages to identify any positive occurrence) and worse on 2. On average, it outperforms competitors in both precision and recall (both micro and macro averaged).
R2Gen (pretrained) R2Gen (fine-tuned) slog
Pathology $Pr$ $Rc$ $F_{1}$ $Pr$ $Rc$ $F_{1}$ $Pr$ $Rc$ $F_{1}$
No Finding 36.00 52.94 42.86 38.35 47.29 42.35 37.65 50.68 43.2
Enl. Cardiomed. 0.51 5.56 0.94 0.00 0.00 0.00 1.61 5.56 2.50
Cardiomegaly 17.00 36.70 23.23 15.66 42.55 22.89 16.64 51.06 25.1
Lung Lesion 0.00 0.00 0.00 0.00 0.00 0.00 20.0 1.67 3.08
Lung Opacity 24.44 8.07 12.13 28.81 12.47 17.41 25.94 15.16 19.14
Edema 44.70 18.97 26.64 40.58 18.01 24.94 40.83 15.76 22.74
Consolidation 0.00 0.00 0.00 4.85 7.04 5.75 12.35 14.08 13.16
Pneumonia 0.00 0.00 0.00 0.00 0.00 0.00 0.0 0.0 0.0
Atelectasis 18.73 22.12 20.28 16.74 15.93 16.33 18.05 27.88 21.91
Pneumothorax 6.82 10.00 8.11 10.64 16.67 12.99 9.62 16.67 12.2
Pleural Effusion 52.68 29.67 37.96 44.55 38.19 41.12 45.05 36.26 40.18
Pleural Other 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Fracture 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Support Devices 18.89 50.0 27.41 18.48 47.73 26.54 18.98 48.86 27.34
Macro avg 14.25 16.72 15.70 15.61 17.56 15.02 17.62 20.26 16.47
Micro avg 25.56 25.77 25.35 25.74 26.57 26.19 25.78 29.13 27.35
5 Related Work
Aligning LLMs.
The standard approach for aligning large language models to human interests is reinforcement learning with human feedback (RLHF) (Ziegler et al., 2020; Ouyang et al., 2022). Several RLHF-based approaches that target medical tasks exist. (Yunxiang et al., 2023; Wang et al., 2023) proposed medical chat models obtained by fine-tuning existing architectures, while (Bazi et al., 2023) introduced a specially designed vision transformer. (Seo et al., 2020) presented a method for improving the performance of an image caption generator with offline human feedback. slog can be viewed as a variant of RLHF that foregoes reinforcement learning in favor of a fully end-to-end fine-tuning strategy, for efficiency. It also differs in aim, in that it optimizes the model’s guidance for a specific human decision making task, rather than for factuality and fairness in general (Ouyang et al., 2022).
Pathological report generation.
Several approaches (Hou et al., 2021; Chen et al., 2020; Wang et al., 2022) have been developed for machine-driven pathological report generation from chest X-ray images using the Mimic-CXR-IV (Johnson et al., 2019) and the Indiana University chest X-ray data sets (Demner-Fushman et al., 2016). (Lovelace and Mortazavi, 2020b) designed a model with a similar objective but they proposed to leverage the CheXpert dataset to enhance the coherence of their model. (Tanida et al., 2023) introduced a region-guided model to generate pathological reports, thus opening the window of interactive human-guided report generation. However, these models are not concerned with optimizing the utility of the generated reports for the follow-up decision making. Our approach builds on top of these methods, enriching them with the ability to incorporate surrogate quality information (we use (Chen et al., 2020) in our evaluation, but any of these models can be adapted for slog).
Other approaches.
LTG is related to explain then predict (ETP) (Camburu et al., 2018; Kumar and Talukdar, 2020; Zhao and Vydiswaran, 2021), a framework for building explainable (Guidotti et al., 2018) models in which a machine first outputs a full-fledged explanation – playing the role of “guidance” – and then derives a prediction from the explanation itself. In LTG, however, the prediction step is carried out by a human expert, and as such it is not differentiable. Also, ETP requires direct supervision on the explanations themselves, which is seldom available. In contrast, slog improves guidance quality using indirect scoring feedback, which is comparatively easier to acquire.
Finally, LTG is not restricted to textual guidance. One option is, for instance, to implement guidance in terms of explanations extracted from (or output by) an underlying image classifier to guide human decision making (Guidotti et al., 2018). From this perspective, LTG is related to explanatory interactive learning (XIL) (Schramowski et al., 2020; Teso et al., 2023), in which the goal is to improve the quality of explanations output by a machine learning model by interactively acquiring corrections to the explanations themselves. The key difference is that LTG focuses on down-stream decision quality and slog supports textual guidance, while XIL aims at more generally improving explanation quality and implementations do no support textual explanations.
6 Conclusion
We introduced learning to guide as an alternative setup for high-stakes hybrid decision making that ensures the human expert is always in the loop, as well as slog, an end-to-end approach for turning pre-trained VLMs into high-quality textual guidance using human feedback. Our results suggest that slog is effective at steering VLMs towards generating more informative guidance, leading to improved accuracy in downstream decisions.
In follow-up work, we plan to extend slog by integrating ideas from active learning (Settles, 2012) to acquire the quality rankings, and to explore connections with explainable AI (Guidotti et al., 2018), explanatory interactive learning (Schramowski et al., 2020; Teso et al., 2023) and skeptical learning (Zeni et al., 2019) to facilitate the identification and correction of potential issues with the generated guidance.
Ethical Statement
The rationale of the LTG paradigm is to improve the quality of decisions for human decision makers in high-stakes applications. By refraining from directly suggesting decisions and focusing on useful guidance, it always leaves the human in charge on the final decision. However, a malicious use of the paradigm could drive humans to make decisions that optimize some undisclosed third-party objective. The requirement for explainability of the guidance combined with expertise of the decision maker should help preventing these potential misuses.